OpenStudy (anonymous):

Two rectangles are similar. If the height of the first rectangle is 3 inches, and the height of the second rectangle is 9 inches, how much longer is the second rectangle's perimeter? 3 times as long 6 times as long 9 times as long 6 inches longer

OpenStudy (anonymous):

6 inches longer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

if you were helped you can click best response to reward your helper ;)

OpenStudy (***[isuru]***):

nope...i think @niels5x9 is mistaken

OpenStudy (anonymous):

true i was wrong i didn't see that last word "perimeter"

OpenStudy (***[isuru]***):

since 2 rectangles are similar... their height to breadth ratio should be equal if the breadth of the first one is "x" and the second one is "y" then 3/x = 9/ y which means y = 3x if the perimeter of the first rectangle is P P = 3 + 3 + x + x = 6 + 2x if the perimeter of the second rectangle is P' P' = 9 + 9 + y + y = 18 + 2y = 18 + 6x = 3 ( 6 + 2x ) = 3 P which means the 3 times the perimeter of the first one :)

OpenStudy (***[isuru]***):

@ltj1234